Write an equivalent expression without using the grouping symbols: -(3x + 5)
Answer by LarissaRichardson(34)

you want to "undo" the brackets
to do this distribute(multiply) the outside of the bracket to the inside
in this case -(3x + 5) = -1 * (3x + 5) = -3x -5